OpenStudy (anonymous):

factorize x^2-34x-573=0

OpenStudy (anonymous):

you need two numbers that multiply to -573 and add up to -34 (i don't think there are any factors that work) i think it's easier to do this using complete the square method. x^2 - 34x = 573 (x^2-34x + 289) = 573 +289 (x-17)^2 = 862 x-17 = + sqrt (862) or - sqrt(862) so x = sqrt(862) + 17 or -sqrt(862) + 17